# Efficient Parallel Algorithms for Inpainting-Based Representations of 4K Images

### v1

@@ Abstract
Homogeneous diffusion inpainting can reconstruct missing image areas with high quality from a sparse subset of known pixels, provided that their location as well as their gray or color values are well optimized. This property is exploited in inpainting-based image compression, which is a promising alternative to classical transform-based codecs such as JPEG and JPEG2000. However, optimizing the inpainting data is a challenging task. Current approaches are either quite slow or do not produce high quality results. As a remedy we propose fast spatial and tonal optimization algorithms for homogeneous diffusion inpainting that efficiently utilize GPU parallelism, with a careful adaptation of some of the most successful numerical concepts. We propose a densification strategy using ideas from error-map dithering combined with a Delaunay triangulation for the spatial optimization. For the tonal optimization we design a domain decomposition solver that solves the corresponding normal equations in a matrix-free fashion and supplement it with a Voronoi-based initialization strategy. With our proposed methods we are able to generate high quality inpainting masks for homogeneous diffusion and optimized tonal values in a runtime that outperforms prior state-of-the-art by a wide margin.
@@ 1 Introduction
Inpainting-based image compression methods [1, 7, 14, 20, 24, 25, 37, 42, 52, 53, 58, 61, 64, 67, 68, 69] can be promising alternatives to classical transform-based codecs. During encoding, they store a small subset of pixels. In the decoding phase, the missing parts of the image are reconstructed by an inpainting operator. Earlier approaches [1, 7, 14, 20, 25, 42, 69] mostly rely on image data at semantic image features such as edge locations. Starting with the work by Galić et al. [23, 24] many later approaches [31, 52, 53, 61] instead carefully optimize the locations of the known data, without any restrictions to semantic image features. In terms of quality, inpainting-based compression methods have been shown to be able to outperform the widely used JPEG [48] and JPEG2000 [66] for images with small to medium amounts of texture [52, 53, 61].
Surprisingly, already with simple linear homogeneous diffusion [33] one can obtain very good inpainting results [8, 11, 12, 15, 28, 44] under the condition that the inpainting data are highly optimized. For piecewise constant images, such as cartoon images, depth maps or flow fields, one can even achieve state-of-the-art results [25, 31, 41, 43] and outperform HEVC [65]. In contrast to nonlinear inpainting methods, homogeneous diffusion inpainting is parameter-free and its discretization yields a sparse system that benefits from its linearity. In Part I of our companion paper [39] we have shown that we can solve this linear system in real-time for 4K color images.
@@ 1.1 Our Contribution
In the first part of our companion paper [39] we have proposed a full multigrid optimized restricted additive Schwarz (ORAS) inpainting solver for homogeneous diffusion inpainting. It achieves real-time decoding with more than $60$ frames per second for 4K color images, by exploiting the potential of highly parallel architectures such as GPUs. Our goal in the second part is to complement our inpainting solver with a fast data optimization framework, in order to also make the encoding for image and video compression with homogeneous diffusion viable for 4K image resolutions. To this end, we propose fast and highly parallel methods to optimize the stored data.
For the spatial optimization we extend the densification approaches by Daropoulos et al. [18], and Chizhov and Weickert [16]. We employ a Delaunay-based densification, which we supplement with a strategy for choosing the initial inpainting mask, based on the analytic mask optimization framework of Belhachmi et al. [8]. Our methods offer state-of-the-art performance. We achieve fast runtimes by efficiently computing the Delaunay triangulation in each densification iteration from an approximation of the Voronoi tessellation, obtained with a jump flood algorithm (JFA) [59].
For our tonal optimization approach, inspired by the nested conjugate gradients (CG) approach by Chizhov and Weickert [16], we solve the corresponding normal equations in a matrix-free fashion. As an outer solver, we adapt a restricted additive Schwarz domain decomposition method [13], which is able to exploit the fact that the influence area of each known pixel is localized in practice. In addition, we devise a computationally inexpensive initialization strategy allowing for a much earlier termination of the tonal optimization. We derive the latter from a theoretically motivated surrogate problem, which also subsumes the initialization from the work of Galić et al. [24].
Since our spatial and tonal optimization methods both rely on the computation of multiple inpaintings, we integrate our fast inpainting solver proposed in the first part of our companion paper [39]. Thanks to all of our contributions, we are able to achieve runtimes of less than a second for the full data optimization of 4K color images and an optimal linear scaling behavior with respect to the image resolution. To the best of our knowledge, this is currently the fastest high-quality data optimization for homogeneous diffusion inpainting.
@@ 1.2.1 Spatial Optimization
Belhachmi et al. [8] developed a framework for mask optimization based on a continuous shape analytic perspective of homogeneous diffusion inpainting, which was later extended by Belhachmi and Jacumin [9] for noisy images. The practically relevant result is that the mask density should be chosen as an increasing function in the absolute value of the Laplacian of the target image. In the discrete setting they suggest applying a dithering method on the absolute value of the smoothed Laplacian image, e.g. Floyd-Steinberg dithering [22], in order to generate a mask that matches the desired density. While this approach is very fast, as it does not require computing any inpaintings, in practice the dithering approximation leads to a limited reconstruction quality [44]. In the current work we aim at a considerably higher quality of the inpaintings, and therefore integrate the Belhachmi approach only as a partial initialization in our algorithm.
@@
Mainberger et al. [44] have proposed a probabilistic sparsification (PS) algorithm that starts with a full mask and iteratively removes pixels from the mask until the desired density is reached. In each iteration, a set of candidate pixels from the mask is chosen and temporarily removed. An error map is computed from the inpainting with the reduced mask. Then a fixed number of the candidates are reintroduced back into the image: the ones with the highest pointwise errors. While this approach can be adapted to different inpainting operators, such as PDE-based ones [44] or linear splines interpolation over triangulations [19], it is computationally fairly expensive. Additionally, since PS starts from a full mask, reaching low densities takes many more iterations than starting from an empty mask and densifying. Last but not least, PS is prone to getting stuck in suboptimal local minima both because it uses only pointwise errors as an oracle, and because the error itself in the initial iterations is a result only of a very localized neighborhood that is not predictive of the influence of mask points in later iterations. These considerations suggest studying densification approaches with an oracle that adapts to the density of the mask in each iteration.
@@
Densification approaches start from an empty mask and iteratively insert new pixels. In practice, this also means that fewer iterations need to be performed compared to sparsification-based methods, since the usual mask densities for inpainting-based compression are below 10%. In compression settings, densification has been successfully applied to constrained data structures such as subdivision trees [21, 24, 52, 61]. In the current work we are interested in the more general case of unconstrained masks. Unconstrained densification strategies have been used successfully for diffusion-based [16, 18, 34] as well as exemplar-based [40] inpainting. The works of Daropoulos et al. [18], Chizhov and Weickert [16], and Jost et al. [34] consider not just pixel-wise errors, but aggregate the error over mask-adaptive neighborhoods based on the Voronoi diagram [18, 34] or the Delaunay triangulation [16], which improves the reconstruction quality compared to a pixel-wise densification. The work by Jost et al. [34] even generalizes the optimization to masks prescribing derivatives and local integrals. Currently, Voronoi- and Delaunay-based densification approaches provide the best quality-to-runtime ratio for mask construction, so we use those as a basis for the mask generation algorithms we develop in the present work.
@@
Even though densification-based methods can perform better than sparsification-based ones, they are still greedy and can thus get stuck in local minima. To alleviate this, a global relocation strategy called nonlocal pixel exchange (NLPE) has been devised by Mainberger et al. [44]. It stochastically chooses a mask pixel and a subset of non-mask pixels with high pointwise errors, and iteratively computes the errors for the reconstructions resulting from relocating the mask pixels to the candidate non-mask pixel locations. The best swap (potentially also none) is kept. While this approach can escape from local minima and achieve significant improvements, it requires a huge number of iterations (i.e. computational resources) to converge. We also note that for linear spline interpolation on triangulations a similar approach has been used to optimize masks [45].
@@
While neural approaches have the advantage of not needing to perform several inpainting steps, model-based optimization is transparent, and the ideas within generalize to a larger class of problems. Therefore, in the current work we rely on non-neural optimization techniques. Thanks to our fast inpainting solver, presented in the companion paper [39], and the efficient optimization algorithms developed in the current work, we are even able to achieve similar runtimes to the method by Schrader et al. [62] while producing higher quality masks.
@@
Most approaches for inpainting-based video compression, such as [3, 4, 54], optimize the inpainting masks for each frame independently. Breuß et al. [12] instead optimize inpainting masks only for key frames, with the optimal control approach of Hoeltgen et al. [28]. For the intermediate frames, they use optical flow to interpolate the masks.
@@ 1.2.2 Tonal Optimization
The system matrix for the normal equations is of the form $\bm{B}^{\top}\bm{B}\in\mathbb{R}^{m\times m}$, where $m$ is the number of mask pixels [30, 44]. The latter is a symmetric positive definite matrix. Such matrices are especially advantageous from a numerical perspective, since algorithms such as the Cholesky decomposition [27] or the conjugate gradient (CG) method [26, 60] can be used to efficiently compute a solution. We note, however, that neither $\bm{B}^{\top}\bm{B}$ nor $\bm{B}$ are sparse. Despite this, the matrix-vector product $\bm{B}\bm{x}$ can be carried out through a sparse matrix-vector product followed by solving a sparse linear system, both of which can be implemented in a matrix-free fashion (i.e. without explicitly forming or storing the matrices).
@@
For smaller images one can compute the matrix $\bm{B}$ explicitly and store it. The columns in this matrix are termed inpainting echoes [44]. Mainberger et al. [44] precompute the echoes and store them, and thus essentially form and store the matrix $\bm{B}$. They note that one could solve the normal equations involving $\bm{B}^{\top}\bm{B}$ with an LU solver [27], but that this is too slow even for $256\times 256$ images, confirming our conclusion that for large images an iterative solver must be used. They employ a modified successive over-relaxation (SOR) solver [60] with an under-relaxation weight and random mask points traversal. In practice, storing $\bm{B}$ is infeasible for larger images due to memory constraints, and recomputing an inpainting echo (i.e. a column of $\bm{B}$) each time a mask value needs to be modified in an iterative approach is also prohibitively expensive. In our method, we use a much more efficient way to evaluate products $\bm{B}\bm{x}$ without ever explicitly forming or storing any matrix.
@@
Homogeneous diffusion inpainting can equivalently be formulated in terms of a linear combination of Green’s functions associated with mask points [5, 30, 32, 38]. The Green’s functions are a kind of spectral counterpart to the inpainting echoes. We will denote the matrix of the Green’s functions, for a specific mask, with $\bm{G}$. Then the corresponding normal equations take the form $\bm{G}^{\top}\bm{G}$ once again [30]. The main difference to the inpainting echoes is that products $\bm{G}\bm{x}$ can be evaluated efficiently and in a matrix-free fashion using the fast orthogonal DCT-II and DCT-III transforms [57]. Hoffman [30] forms $\bm{G}^{\top}\bm{G}$ explicitly and uses a Cholesky decomposition on it. He demonstrates that for a low number of mask points the Choleksy approach applied to the Green’s functions formulation can be more efficient than conjugate gradients applied to the normal equations of the classical formulation. Since we are dealing with many more mask points due to considering $4K$ images (in Hoffman’s case $4\%$ of a $256\times 256$ image), it is clear that the above approach is impractical in our setting.
For 1D signals, Plonka et al. [56] proposed a method, based on the Green’s functions approach, that jointly optimizes the locations of the mask pixels as well as their tonal values. As this approach only considers 1D signals it is not applicable in our setting and since it uses Green’s functions would suffer from similar restrictions as the approach by Hoffman.
@@
A potential remedy of the memory constraint problem was proposed by Chizhov and Weickert [16] for their finite elements framework. They use a CG solver [26, 60] on the normal equations, but instead of explicitly computing the system matrix they evaluate the matrix-vector products $\bm{B}\bm{x}$ and $\bm{B}^{\top}\bm{x}$ in an efficient manner. They rewrite $\bm{B}=\tilde{\bm{A}}^{-1}\tilde{\bm{C}}$, where $\tilde{\bm{C}}$ is sparse, and where $\tilde{\bm{A}}$ is sparse and symmetric positive definite. Therefore, they use a conjugate gradients solver whenever they have to compute $\tilde{\bm{A}}^{-1}\bm{y}$ allowing them to exploit the sparsity of the problem. With this so-called nested CG approach, they achieve a memory and runtime efficient tonal optimization. We use a similar nested strategy in our framework, where the nested iterations are replaced by an adapted version of the fast inpainting algorithm from our companion paper [39].
@@
Since the influence zone of a single mask pixel is mostly limited for many inpainting operators, localization strategies are an option to accelerate tonal optimization. This works especially well for perfectly localized inpainting operators such as Shepard interpolation with truncated Gaussians [49], 1-D linear interpolation [51], or smoothed particle hydrodynamics [18]. For more global operators, there are also methods that artificially limit the influence by subdivision trees [53] or segmentation [31, 36]. Similarly, one can also restrict the support of inpainting echoes in homogeneous diffusion [30]. This essentially constructs an approximation $\tilde{\bm{B}}$ of $\bm{B}$ that is not fully dense anymore. Nevertheless, the number of restricted echoes which have to be computed and stored is still very large for 4K images.

### v2

# Efficient Parallel Data Optimization for

@@ Abstract
Homogeneous diffusion inpainting can reconstruct missing image areas with high quality from a sparse subset of known pixels, provided that their location as well as their gray or color values are well optimized. This property is exploited in inpainting-based image compression, which is a promising alternative to classical transform-based codecs such as JPEG and JPEG2000. However, optimizing the inpainting data is a challenging task. Current approaches are either fairly slow or do not produce high quality results. As a remedy we propose fast spatial and tonal optimization algorithms for homogeneous diffusion inpainting that efficiently utilize GPU parallelism, with a careful adaptation of some of the most successful numerical concepts. We propose a densification strategy using ideas from error-map dithering combined with a Delaunay triangulation for the spatial optimization. For the tonal optimization we design a domain decomposition solver that solves the corresponding normal equations in a matrix-free fashion and supplement it with a Voronoi-based initialization strategy. With our proposed methods we are able to generate high quality inpainting masks for homogeneous diffusion and optimized tonal values in a runtime that outperforms prior state-of-the-art by a wide margin.
@@ 1 Introduction
Inpainting-based image compression methods [1, 7, 19, 25, 32, 34, 47, 52, 62, 67, 70, 73, 84, 85, 86, 35] can be promising alternatives to classical transform-based codecs. During encoding, they store a small subset of pixels. In the decoding phase, the missing parts of the image are reconstructed from the stored pixels by an inpainting operator. Earlier approaches [1, 7, 19, 25, 34, 52, 86] mostly rely on image data at semantic image features such as edge locations. Many later approaches [32, 42, 62, 70] instead carefully optimize the locations of the known data, without any restrictions to semantic image features. In terms of quality, inpainting-based compression methods have been shown to be able to outperform the widely used JPEG [59] and JPEG2000 [77] for images with small to medium amounts of texture [62, 70].
Surprisingly, already with simple linear homogeneous diffusion [44] one can obtain very good inpainting results [8, 13, 16, 20, 39, 54] under the condition that the inpainting data are highly optimized. For piecewise constant images, such as cartoon images, depth maps or flow fields, one can even achieve state-of-the-art results [34, 42, 51, 53] and outperform HEVC [76]. In contrast to nonlinear inpainting methods, homogeneous diffusion inpainting is parameter-free and its discretization yields a sparse system that benefits from its linearity. In [49] it has been shown that we can solve this linear system in real-time for 4K color images.
@@ 1.1 Our Contribution
The goal of our paper is to develop a state-of-the-art data optimization framework for homogeneous diffusion inpainting that complements the fast inpainting solver from [49]. To this end, we devise fast and highly parallel methods to optimize the data to be stored. This allows us to construct the first viable data optimization approach for sparse 4K image inpainting in high quality that also exploits tonal optimization.
For the spatial optimization we extend the densification approaches by Daropoulos et al. [23], and Chizhov and Weickert [21]. We employ a Delaunay-based densification, which we supplement with a strategy for choosing the initial inpainting mask, based on the analytic mask optimization framework of Belhachmi et al. [8]. Our methods offer state-of-the-art performance. We achieve fast runtimes by efficiently computing the Delaunay triangulation in each densification iteration from an approximation of the Voronoi tessellation, obtained with the jump flood algorithm (JFA) [68].
For our tonal optimization approach, inspired by the nested conjugate gradients (CG) approach by Chizhov and Weickert [21], we solve the corresponding normal equations in a matrix-free fashion. Instead of relying on CG, we construct a novel highly parallel solver for tonal optimization adapted to our sparse inpainting problem. It carefully adapts and integrates ideas from multigrid and domain decomposition, see Figure 6. As an outer solver, we adapt a restricted additive Schwarz domain decomposition method [18], which is able to exploit the fact that the influence area of each known pixel is localized in practice. In addition, we devise a computationally inexpensive initialization strategy allowing for a much earlier termination of the tonal optimization. We derive the latter from a theoretically motivated surrogate problem, which also subsumes the initialization from the work of Galić et al. [32].
Since our spatial and tonal optimization methods both require the computation of multiple inpainting solutions, we use the multigrid ORAS inpainting solver proposed in [49]. This is to the best of our knowledge the fastest homogeneous diffusion inpainting solver for parallel architectures such as GPUs. Thanks to all of our contributions, we are able to achieve runtimes of less than a second for the full data optimization of 4K color images with an optimal linear scaling behavior with respect to the image resolution. Our tonal optimization is currently the fastest method for homogeneous diffusion inpainting, taking less than half a second for a 4K RGB image. Our spatial optimization is the second fastest available mask optimization approach, also taking less than half a second for the construction of outstanding quality masks for a 4K RGB image. The only faster spatial optimization strategy is the method by Belhachmi et al. [8], however, it results in inpaintings that can have a PSNR of up to $6$ dB lower compared to our approach, see Figure 10.
@@ 1.2.1 Spatial Optimization
Belhachmi et al. [8] developed a framework for mask optimization based on a continuous shape analytic perspective of homogeneous diffusion inpainting, which was later extended by Belhachmi and Jacumin [10, 11] for noisy images. The practically relevant result is that the mask density should be chosen as an increasing function in the absolute value of the Laplacian of the target image. In the discrete setting they suggest applying a dithering method to the absolute value of the smoothed Laplacian image, e.g. Floyd-Steinberg dithering [31], in order to generate a mask that matches the desired density. While this approach is very fast, as it does not require computing any inpaintings, in practice the dithering approximation leads to a limited reconstruction quality [54]. In the current work we aim at a considerably higher quality of the inpaintings, and therefore integrate the Belhachmi approach only as a partial initialization in our algorithm.
@@
Mainberger et al. [54] have proposed a probabilistic sparsification (PS) algorithm that starts with a full mask and iteratively removes pixels from the mask until the desired density is reached. In each iteration, a set of candidate pixels from the mask is chosen and temporarily removed. An error map is computed from the inpainting with the reduced mask. Then a fixed number of the candidates are reintroduced back into the image: the ones with the highest pointwise errors. While this approach can be easily adapted to different inpainting operators, such as PDE-based ones [54] or linear splines interpolation over triangulations [24], it is computationally fairly expensive. Additionally, since PS starts from a full mask, reaching low densities takes many more iterations than starting from an empty mask and densifying. Last but not least, PS is prone to getting stuck in suboptimal local minima both because it uses only pointwise errors as an oracle, and because the error itself in the initial iterations is a result only of a very localized neighborhood that is not predictive of the influence of mask points in later iterations. These considerations suggest studying densification approaches with an oracle that adapts to the density of the mask in each iteration.
@@
Densification approaches start from an empty mask and iteratively insert new pixels. In practice, this also means that fewer iterations need to be performed compared to sparsification-based methods, since the usual mask densities for inpainting-based compression are below 10%. In compression settings, densification has been successfully applied to constrained data structures such as subdivision trees [27, 32, 62, 70]. In the current work we are interested in the more general case of unconstrained masks. Unconstrained densification strategies have been used successfully for diffusion-based [9, 21, 23, 45, 80] and exemplar-based [50] inpainting, as well as as linear interpolation on Delaunay triangulations [3, 30]. The works of Daropoulos et al. [23], Chizhov and Weickert [21], and Jost et al. [45] consider not just pixel-wise errors, but aggregate the error over mask-adaptive neighborhoods based on the Voronoi diagram [23, 45] or the Delaunay triangulation [21], which improves the reconstruction quality compared to a pixel-wise densification. The work by Jost et al. [45] even generalizes the optimization to masks prescribing derivatives and local integrals. Currently, Voronoi- and Delaunay-based densification approaches provide the best quality-to-runtime ratio for mask construction, so we use those as a basis for the mask generation algorithms we develop in the present work.
@@
Even though densification-based methods can perform better than sparsification-based ones, they are still greedy and can thus get stuck in local minima. To alleviate this, a global relocation strategy called nonlocal pixel exchange (NLPE) has been devised by Mainberger et al. [54]. It stochastically chooses a mask pixel and a subset of non-mask pixels with high pointwise errors, and iteratively computes the errors for the reconstructions resulting from relocating the mask pixels to the candidate non-mask pixel locations. The best swap (potentially none) is kept. While this approach can escape from local minima and achieve significant improvements, it requires a huge number of iterations (i.e. computational resources) to converge. We also note that for linear spline interpolation on triangulations a similar approach has been used to optimize masks [55].
@@
While neural approaches have the advantage of not needing to perform several inpainting steps, model-based optimization is transparent, and the ideas within generalize to a larger class of problems. Therefore, in the current work we rely on non-neural optimization techniques. Thanks to using the fast inpainting solver, proposed in [49], and the efficient optimization algorithms developed in the current work, we are even able to achieve similar runtimes to the method by Schrader et al. [71] with much lower memory usage, while producing higher quality masks.
@@
Most approaches for inpainting-based video compression, such as [5, 63], optimize the inpainting masks for each frame independently. Breuß et al. [16] instead optimize inpainting masks only for key frames, with the optimal control approach of Hoeltgen et al. [39]. For the intermediate frames, they use optical flow to interpolate the masks.
@@ 1.2.2 Tonal Optimization
The system matrix for the normal equations is of the form $\bm{B}^{\top}\bm{B}\in\mathbb{R}^{m\times m}$, where $m$ is the number of mask pixels [41, 54]. The latter is a symmetric positive definite matrix. Such matrices are advantageous from a numerical perspective, since algorithms such as the Cholesky decomposition [38] or the conjugate gradient (CG) method [37, 69] can be used to efficiently compute a solution. We note, however, that neither $\bm{B}^{\top}\bm{B}$ nor $\bm{B}$ are sparse. Despite this, the matrix-vector product $\bm{B}\bm{x}$ can be carried out through a sparse matrix-vector product followed by solving a sparse linear system, both of which can be implemented in a matrix-free fashion (i.e. without explicitly forming or storing the matrices).
@@
For smaller images one can compute the matrix $\bm{B}$ explicitly and store it. The columns in this matrix are termed inpainting echoes [54]. Mainberger et al. [54] precompute the echoes and store them, and thus essentially form and store the matrix $\bm{B}$. They note that one could solve the normal equations involving $\bm{B}^{\top}\bm{B}$ with an LU solver [38], but that this is too slow even for $256\times 256$ images. This supports our statement that for large images an iterative solver has to be used. They employ a modified successive over-relaxation (SOR) solver [69] with an under-relaxation weight and random mask points traversal. In practice, storing $\bm{B}$ is infeasible for larger images due to memory constraints, and recomputing an inpainting echo (i.e. a column of $\bm{B}$) each time a mask value needs to be modified in an iterative approach is also prohibitively expensive. In our method, we use a much more efficient way to evaluate products $\bm{B}\bm{x}$ without ever explicitly forming or storing any matrix.
@@
Homogeneous diffusion inpainting can equivalently be formulated in terms of a linear combination of Green’s functions associated with mask points [41, 43, 48]. The Green’s functions are a kind of spectral counterpart to the inpainting echoes. We will denote the matrix of the Green’s functions, for a specific mask, with $\bm{G}$. Then the corresponding normal equations take the form $\bm{G}^{\top}\bm{G}$ once again [41]. The main difference to the inpainting echoes is that products $\bm{G}\bm{x}$ can be evaluated efficiently and in a matrix-free fashion using the fast orthogonal DCT-II and DCT-III transforms [66]. Hoffman [41] forms $\bm{G}^{\top}\bm{G}$ explicitly and uses a Cholesky decomposition on it. He demonstrates that for a low number of mask points the Choleksy approach applied to the Green’s functions formulation can be more efficient than conjugate gradients applied to the normal equations of the classical formulation. Since we are dealing with many more mask points due to considering $4K$ images (in Hoffman’s case this was just $4\%$ of a $256\times 256$ image), it is clear that the above approach is impractical in our setting.
For 1D signals, Plonka et al. [65] have proposed a method, based on the Green’s functions approach, that jointly optimizes the locations of the mask pixels as well as their tonal values. As this approach only considers 1D signals it is not applicable in our setting, and since it uses Green’s functions it suffers from similar restrictions as the approach by Hoffman.
@@
A potential remedy of the memory constraint problem was proposed by Chizhov and Weickert [21] for their finite elements framework. They use a CG solver [37, 69] on the normal equations, but instead of explicitly computing the system matrix they evaluate the matrix-vector products $\bm{B}\bm{x}$ and $\bm{B}^{\top}\bm{x}$ in an efficient manner. They rewrite $\bm{B}=\tilde{\bm{A}}^{-1}\tilde{\bm{C}}$, where $\tilde{\bm{C}}$ is sparse, and where $\tilde{\bm{A}}$ is sparse and symmetric positive definite. Therefore, they use a conjugate gradients solver whenever they have to compute $\tilde{\bm{A}}^{-1}\bm{y}$ allowing them to exploit the sparsity of the problem. With this so-called nested CG approach, they achieve a memory and runtime efficient tonal optimization. In our approach, we use a similar nested strategy in our framework, where the nested iterations are replaced by an adapted version of the fast inpainting algorithm from [49].
@@
Since the influence zone of a single mask pixel is mostly limited for many inpainting operators, localization strategies are an option to accelerate tonal optimization. This works especially well for perfectly localized inpainting operators such as Shepard interpolation with truncated Gaussians [2, 60], or smoothed particle hydrodynamics [23]. For less local operators, there are also methods that artificially limit the influence with a segmentation [42, 46]. Similarly, one can also restrict the support of inpainting echoes in homogeneous diffusion [41]. This essentially constructs an approximation $\tilde{\bm{B}}$ of $\bm{B}$ that is not fully dense anymore. Nevertheless, the number of restricted echoes which have to be computed and stored is still very large for 4K images.